### v1

# Planning Anything with Rigor: General-Purpose Zero-Shot Planning with LLM-based Formalized Programming

Abstract
While large language models (LLMs) have recently demonstrated strong potential in solving planning problems, there is a trade-off between flexibility and complexity. LLMs, as zero-shot planners themselves, are still not capable of directly generating valid plans for complex planning problems such as multi-constraint or long-horizon tasks. On the other hand, many frameworks aiming to solve complex planning problems often rely on task-specific preparatory efforts, such as task-specific in-context examples and pre-defined critics/verifiers, which limits their cross-task generalization capability. In this paper, we tackle these challenges by observing that the core of many planning problems lies in optimization problems: searching for the optimal solution (best plan) with goals subject to constraints (preconditions and effects of decisions). With LLMs’ commonsense, reasoning, and programming capabilities, this opens up the possibilities of a universal LLM-based approach to planning problems. Inspired by this observation, we propose LLMFP, a general-purpose framework that leverages LLMs to capture key information from planning problems and formally formulate and solve them as optimization problems from scratch, with no task-specific examples needed. We apply LLMFP to 9 planning problems, ranging from multi-constraint decision making to multi-step planning problems, and demonstrate that LLMFP achieves on average 83.7% and 86.8% optimal rate across 9 tasks for GPT-4o and Claude 3.5 Sonnet, significantly outperforming the best baseline (direct planning with OpenAI o1-preview) with 37.6% and 40.7% improvements. We also validate components of LLMFP with ablation experiments and analyzed the underlying success and failure reasons.
1 Introduction
Figure 1: An overview of LLMFP and how it is applied to a coffee supply chain example. All sections in yellow are inputs, and all sections in blue are steps accomplished by LLMs. With task description, background information, and query as inputs, LLMFP defines the goal, decision variables, and constraints of this optimization problem, identifies all necessary variables and summarizes their key information into a JSON representation, generates codes to solve the optimization problem, executes the codes and formats the execution results, and performs self-assessment for each step.
Making complex plans subject to multiple constraints is a time- and labor-intensive process, but is critical in many aspects of our lives such as work arrangement, business management, logistics, and robotics. In the past, people used domain-specific tools and languages to make specific plans in their areas, which often required a steep learning curve and were hard to adapt to other domains. When large language models (LLMs) emerge with their versatile capabilities such as language understanding, reasoning, and tool-using, using LLMs for planning has gained significant traction.
For such planning systems to be deployed in complex, real-world applications, two desirable properties need to be satisfied: 1). Zero-shot flexibility: Unlike many experimental settings where planning tasks usually come with labeled datasets, it is very challenging to request such datasets from users in many realistic settings. Ideally, a flexible planning system should be able to conduct planning with only a task description provided by users, and nothing else. 2). High performance on complex tasks: Realistic planning problems usually require multi-step, long-horizon solutions, with many explicit and implicit constraints.
However, there is a trade-off between flexibility and task complexity, and thus existing LLM-based planning systems are typically unable to achieve both simultaneously. On one hand, planning systems capable of performing zero-shot planning, utilizing the abundant knowledge and generalization capabilities in LLMs, in many successful applications can only solve single-objective tasks such as household chores, with step-by-step interactive planning and grounding (Huang et al., 2022a; Ahn et al., 2022; Huang et al., 2022b). For complex, multi-constraint, and long-horizon tasks that involve iterative trials and errors even for humans, LLMs still do not have the capabilities to generate valid plans by themselves (Kambhampati et al., 2024). On the other hand, recent research efforts to empower LLM-based planners to solve complex tasks are often based on well-designed task-specific in-context examples and extensive task-specific pre-defined efforts (Liu et al., 2023; Xie et al., 2023; Li et al., 2023; Song et al., 2023; Gundawar et al., 2024), impacting their zero-shot flexibility. In short, few existing planning systems can flexibly resolve generic complex tasks with only task descriptions in natural language. Hence we ask: Can we build a universal LLM-based planning system that can solve complex planning problems without task-specific efforts?
In this paper, we observe that, although planning problems come with drastic variations, many of them can be recast as constrained optimization problems — The optimization problems aim to find the optimal solution, which is equivalent to locating the best plan that satisfies the goal for planning problems; the decision’s precondition and effect are equivalent to constraints of optimization problems. Furthermore, although solving complex planning tasks is generally beyond the capabilities of LLMs, converting any planning tasks into optimization problems is a much more tractable problem, and can be within the zero-shot capabilities of LLMs.
Motivated by this, we propose LLM-Based Formalized Programming (LLMFP, illustrated in Fig. 1), a general-purpose zero-shot planning framework that leverages LLM’s strong common reasoning, and programming capabilities to encode planning problems into optimization problems without any task-specific examples or designs, combined with a formal planner to solve the optimization problem. LLMFP takes in natural language domain description, natural language query under this domain, and available background information or APIs as inputs, and solves the planning problem in five steps. First, LLMFP prompts LLMs to reason and propose the goal, decision variables, and key constraints necessary for the task. Second, based on the response, LLMFP asks LLMs to formulate a representation that includes all variables needed to construct and their detailed information and requirements. Third, with the representation, LLMs write codes to formally encode the problem into an optimization problem. Fourth, LLMFP executes the generated codes, converts the execution results into plans. Finally, LLMFP performs overall self-assessment and automatic modification to fix the broken parts of the previous steps. Currently, LLMFP uses the satisfiability modulo theory (SMT) to encode the optimization problems but can be adapted to any planners or solvers by updating the requirements and representation format in the prompts.
We evaluate our framework with 9 diverse planning problems, ranging from single-step supply chain problem to multi-step robot block stacking and moving. Experiment results demonstrate that LLMFP achieves strong performance across all tasks with an average of 83.7% and 86.8% optimal rates for GPT-4o and Claude 3.5 Sonnet, which greatly outperforms the baselines, including direct plan generation with OpenAI o1-preview. We conduct ablation experiments to validate the key components of our framework and investigate the underlying reasons why our framework is more effective than baselines. In addition, although our framework does not require task-specific examples, we show the ease of adding task-specific examples to one stage of LLMFP, and how it could help to clarify unclear queries and therefore can further improve the performance within the same domain.
In summary, our key contributions are:
  • We offer a novel perspective on using LLMs to solve planning problems by rigorously constructing optimization problems from scratch, alike how human experts use optimization tools for planning.
  • We propose LLMFP, a general-purpose planning framework with zero-shot generalization capability. To our knowledge, LLMFP is the first to enable LLMs to build and solve diverse types of planning problems as optimization problems with no task-specific examples or external critics.
  • LLMFP notably achieves 83.7% and 86.8% optimal rates for GPT-4o and Claude 3.5 Sonnet, outperforming the best baseline (direct planning with OpenAI o1-preview) by 37.6% and 40.7%. We examine the effectiveness of our framework and analyze the success and failure reasons.
2 Related Works
2.1 LLMs for Planning
The remarkable capabilities of LLMs in reasoning (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2022; 2024) and tool-use (Qin et al., 2023; Schick et al., 2024) brings up interests of utilizing LLMs to solve planning problems. Based on LLMs commonsense and zero-shot generalization capability, many works capable of performing zero-shot planning are proposed (Huang et al., 2022a; Ahn et al., 2022; Huang et al., 2022b). However, their planning scenarios are limited to simple single-objective tasks such as household cleaning and they often require step-by-step interactive planning with grounding. To improve LLMs reasoning and planning capabilities for more complex problems, chain-of-thought (CoT) prompting appends reasoning steps before the answer to encourage LLMs to perform step-by-step reasoning (Wei et al., 2022). Recent works also propose to aid the LLM planning processes with external tools (Liu et al., 2023; Guan et al., 2023; Chen et al., 2023; Li et al., 2023; Gundawar et al., 2024; Hao et al., 2024; Chen et al., 2024). For example, (Liu et al., 2023; Xie et al., 2023; Gundawar et al., 2024) leverages LLMs as translators to convert problems into fixed formats and inputting them to external planners, (Li et al., 2023) prompts LLM to add short codes to existing optimization codes of certain domain to account for follow-up what-if questions, and (Gundawar et al., 2024; Chen et al., 2024) empowers LLMs to iteratively refine plans or prompts based on feedback from external task-specific critics/verifiers/humans. However, to achieve strong performance, these methods are often based on extensive task-specific pre-defined efforts. For example, CoT depends on task-specific examples to achieve strong performance, domain files written in the planning domain definition language (PDDL) (Aeronautiques et al., 1998; Haslum et al., 2019) are required for (Liu et al., 2023), codes that encode domains into a mixed-integer linear program (MILP) are necessary for (Li et al., 2023), and external constraint critics are needed for (Gundawar et al., 2024). These requirements limit the generalization capability of these works to new domains.
2.2 LLM + Solver
As existing LLMs still do not have the capability to perform long-horizon reasoning for complex tasks (Achiam et al., 2023; Valmeekam et al., 2022; 2023; Kambhampati et al., ), many works propose to take advantages of both LLMs and external planners or solvers by combining them for reasoning or planning. (Wu et al., 2022; He-Yueya et al., 2023; Pan et al., 2023; Ye et al., 2024) combines LLM with symbolic solvers to solve logical reasoning problems. LLMFP differs from them in that we aim to solve complex planning problems, which could include implicit constraints not clearly described in the task description and could be sequentially long-horizon tasks with defined actions. Instead, most logical reasoning problems are single-step satisfiction problems and clearly specify all constraints in question. In addition, LLMFP proposes a general approach, which does not require task-specific examples or task-specific efforts. (Li et al., 2023) teaches LLMs to add constraints to existing MILP code snippets of planning problems. (Li et al., 2024) asks the developer to express planning problems into automaton and guide the LLMs for planning based on it. (Liu et al., 2023; Guan et al., 2023; Zhou et al., 2024; Xie et al., 2023) leverages PDDL planner to aid the planning processes. Except for the natural language task description, they require human efforts to design solver-related specifications and task-specific examples, which is not needed for LLMFP.
3 LLMFP
LLMFP aims to deliver a plan that can resolve generic planning problems. For example, consider a coffee supply chain problem, where a coffee company sources beans from three suppliers with fixed capacity, roasts them at two facilities into either dark or light coffee and ships the roasted coffee to three retail locations to fulfill their demands. Then a planning problem involves accomplishing the task at the cheapest cost.
To achieve this, LLMFP takes the following inputs from users, as shown in Figure 1 (top panels).
  • Natural Language Task Description. A natural language description that details the problem settings and the planning objective, such as the above description of the coffee problem.
  • Background Information & API. A list of background information about the tasks as well as information on APIs that the planner can use. An example of the background information for the coffee task is the variables containing specific numbers of supplier capacities, cafe demands, and costs for shipping and roasting.
  • User Query. The question that either describes the detailed initial and/or goal states or adds/modifies existing requirements of the tasks. In the coffee planning task, one example query is ‘What would be the outcome if cafe1 experienced a 23% rise in demand’.
Example inputs for all 9 tasks can be found in Appendix A.5. Note that LLMFP does not require any task-specific examples from the users. Considering the diversity of user requests, LLMFP needs to accommodate a large variety of domains, planning problem setups, user queries, constraints, and complexity levels, which poses a great challenge.
3.1 Overview
Emulating how humans construct and write codes to solve optimization problems, LLMFP solves the planning problems via the following steps, as shown in Fig. 1. ❶ Definer: LLMFP first prompts an LLM to define the problem by analyzing and proposing the optimization goal, decision variables, and constraints of the optimization problem (Sec. 3.2). ❷ Formulator: LLMFP asks LLM to think about the necessary variables and steps to build when programming, and formulate a representation to summarize all key information of these variables (Sec. 3.3). ❸ Code Generator: Given this representation, LLMs generate codes that initialize all necessary variables, assert constraints, and add goals (Sec. 3.4). ❹ Result Formatter: After LLMFP executes the generated codes, it prompts LLMs to convert the execution result into a fixed format and provide a short assessment of the execution results (Sec. 3.5). ❺ Self Assessment and Modification: LLMFP assesses each step based on the execution result, and modifies the first incorrect step (Sec. 3.6). The generated plan is delivered when it passes self-assessments of all steps. Please refer to Appendix A.6 and A.7 for example outputs and prompts of all steps in LLMFP.
3.2 Definer
The first step of building an optimization problem is to identify the goal, decision variables, and constraints of the problem. This is accomplished by prompting the LLM to express in a natural language format (See Figure 1 for an example). The prompt includes ❶ all the user-supplied task information; ❷ a description of what goal, decision variables, and constraints mean; and ❸ an instruction to output the aforementioned information. The detailed prompt is listed in Appendix A.7.2.
While generating the goals and decision variables are straightforward, generating the constraints is challenging, because certain constraints are not explicitly stated and can only be inferred by commonsense reasoning. We refer to these as the implicit constraints.
For example, in the coffee supply chain task example, the implicit constraints include ‘the roasted coffee in each roastery does not exceed the beans it receives’, ‘the shipped coffee from each roastery does not exceed the coffee it roasts’, and importantly but easily overlooked, ‘all numbers of shipped and roasted beans and coffee need to be non-negative integer’.
To facilitate uncovering the implicit constraints, we include in the prompts (under the description of constraints) a three-step instruction to derive the constraints: ❶ Identify all decision variables in this task, ❷ for each pair of decision variables, consider relations (explicit, implicit, underlying assumption, unmentioned commonsense) between them to make sure all variables are connected, and ❸ provide a constraint reasoning first before answering. This effectively helps LLMs to better identify implicit constraints for multi-constraint planning problems. Since for multi-step planning problems the task description needs to explicitly define the preconditions and effects of each action, there will be no implicit constraint so this step is omitted.
3.3 Formulator
Figure 2: Examples of diverse variables and steps appeared in different tasks. Examples 1 and 2 are from single-step multi-constraint problems, and Examples 3 and 4 are from multi-step problems.
Although from the last step, we have defined the optimization goal, decision variables, and constraints, more efforts are needed to integrate these into executable and correct code snippets. For example, the decision variables may be continuous, binary, or integer; they may need to be arranged into data structures such as lists and dictionaries; they may have fixed values or belong to flexible ranges, etc. All this information and requirements need further declarations to better generate codes. Fig. 2 shows examples of diverse types of variables.
Single-Step Multi-Constraint Problem To foster LLM to better formulate the variables, we prompt LLM to utilize a fixed-format JSON representation to summarize all variables and their key information. As shown in Fig. 1 and Fig. 2, for each variable, the representation uses 6 fields to summarize the information related to this variable. The name field indicates the variable name. Since we are using SMT as the optimization solver. the SMT_variable field indicates whether the variable is an SMT variable or a normal variable. SMT variables are different from normal variables in that they don’t hold specific values upfront, rather, they are symbolic variables to represent unknown values. The number_of_variables field represents the length of the variable. For example, the total cost in the coffee supply chain task has a length of 1, but the number of coffee beans shipped from supplier to roasteries would be a list of length 6 since there are 3 suppliers and 2 roasteries. The data_source field denotes the dependencies of the variable. For example, since the total cost of the coffee task is the summation of shipping cost from suppliers to roasteries, roasting cost, and shipping cost from roasteries to cafes, they are the data source of the total cost. The value field further specifies the value of the variable. It could either be a real number or list, a string describing keys and values when the data structure of the variable is a dictionary, or it could be a description of operations to do with the data source. In the coffee example, the value is a string “sum of all costs” for the total cost variable. The specific_requirement field is where we point to the constraints or goals related to this variable. For example, the specific requirement for total cost is “minimize”, and for the number of coffee beans shipped from supplier to roasteries, it could be “non-negative integer, summation for each supplier greater than capacity”. With this intermediate step between Definer and Code Generator, LLMFP is capable of obtaining a more detailed, well-formulated, and overall coding plan. We teach LLM to formulate variables into this representation by including two simple examples in the prompt. Note these two examples are not task-specific examples of any of our testing tasks and we do not modify these two examples across tasks.
Multi-Step Planning Problem In addition to including information of variables in the JSON representation, more information is needed for multi-step planning problems. Since predicates, actions, and their preconditions and effects are involved in multi-step problems, we separate the JSON representation for multi-step problems into five stages: objects, predicates, actions, update, goal and each stage includes 6-field representations described above. The objects stage declares all objects in the scenario. The predicates stage defines the predicates, which represent the properties of objects and the relationships between them. The actions stage initializes variables to represent all actions. The update stage adds assertions to existing action variables to account for the preconditions and effects of actions. The goal stage adds constraint to existing predicate variables to encode the goal. As shown in Fig. 2, Example 3 shows a variable in predicates stage, which initializes a dictionary with keys (block, block, timestep) to represent whether a block is on another block at a certain timestep. Example 4 shows a step in update stage, which collects all defined actions in actions stage and asserts that only one action can be true for a timestep. For multi-step planning tasks, we replace the examples with a multi-step task example, and similarly, it is not a task-specific example and we do not modify it across tasks.
To generate the above information, the Formulator’s prompt includes ❶ A brief instruction for the Formulator, ❷ Example input-output pairs of Formulator as demonstrations, and ❸ The user-provided task information and Definer’s output. The detailed prompt is listed in Appendix A.7.2. Note that although example input-output pairs are used, they are task-agnostic examples fixed for all the planning tasks. No task-specific examples are needed.
3.4 Code Generator
With the representation generated from the Formulator, now we have all the information needed to build an optimization problem with codes. In the Code Generator’s prompt, we explain the meanings of different stages and fields in the JSON representation and ask LLMs to follow Python and Z3 SMT syntax (De Moura & Bjørner, 2008). By including user-provided task information and results from Definer and Formulator, with no examples, LLM could reliably generate reasonable, executable, and correct Python codes. Then, LLMFP executes the codes and returns to re-generation if there are runtime errors. We set our maximum re-generation times to be 5.
3.5 Result Formatter
Since the variable names are decided by LLMs and have chances to be very different across queries, after code generation, we use a Result Formatter to ask LLM to convert the execution result to a fixed output format. For example, the output for the coffee task would be a JSON with the following fields: ❶ number of coffee beans shipped from each supplier to each roastery, ❷ the number of light and dark coffee roasted in each roastery, and ❸ the number of light and dark coffee shipped from each roastery to each cafe. After filling in this result, we prompt the LLM to provide a brief evaluation of the result based on whether the result achieves the goal, satisfies constraints, and makes sense in common sense. Taking commonsense into consideration is important because sometimes if a necessary constraint is missing from the Definer step, it could result in unreasonable execution result thus could be unrealistic in commonsense. For example, for the coffee task, if the Definer does not include the non-negative constraint, to minimize the cost, the solver could propose negative units of shipped coffee. Detecting these unrealistic plans is helpful in the Self Assess & Modification step.
3.6 Self Assess & Modification
After we have the execution result and evaluation, LLMFP perform self-assessment to reason about the correctness and provide a rating for the Definer, Formulator, and Code Generator. If the assessment marks all three steps to be correct, this plan will be delivered as the final output of our framework. Otherwise, the assessment will reason about how to modify this step, and provide a modification by itself. This modification will replace the output of the incorrect step and LLMFP will loop back to continue the next steps from there again. That is, if the self Assessor thinks the output of the Formulator is incorrect, it will generate a JSON representation by itself, and the framework will use this modified representation to enter Code Generator again. We set the maximum number of loops to be 5.
3.7 Choice of Solver
As a framework that proposes to formulate and solve planning problems as optimization problems, LLMFP could be adapted to use any planner or solver by modifying the requirements in prompt to follow the syntax of new solvers. In this work, we compare the SMT solver with popular PDDL and MILP solvers and choose SMT solver with following reasons: SMT allows explicit goal and constraint assertion from scratch, which could be used to solve both single-step multi-constraint problems and multi-step problems, while PDDL solvers require a PDDL domain file and a PDDL problem file with strictly fixed formats, which limits its capability to solve non-PDDL problems. SMT is complete and sound, that is, it guarantees to find the optimal plan, while PDDL planners are not guaranteed to be complete. In addition, for all optimization solvers like SMT and MILP, the processes of building optimization problems for different optimization solvers or planners are the same: defining the goal, constraints, and decision variables, and writing codes to encode relationships between decision variables. Thus, utilizing any optimization planner has a similar process. We show how easily our framework could adapt to use MILP by including prompt differences and output examples in Appendix A.7.3. We selected SMT over MILP because the SMT Z3 solver is a publicly available package and is more accessible to all users than the Gurobi MILP solver, which requires licenses and limits the number of devices per license.
4 Experimental Results
4.1 Domains
We test on 9 planning problems, which includes 5 multi-constraint decision making tasks,Coffee, Workforce, Facility, Task Allocation, and Warehouse, and 4 multi-step tasks, Blocksworld, Mystery Blocksworld, Movie, and Gripper (Li et al., 2023; Valmeekam et al., 2024; Stein & Koller, 2023). Detailed task descriptions are included in Appendix A.1. The queries are either what-if questions that change/add constraints to the existing scenarios or different task initial and goal conditions. Task inputs including example queries are given in Appendix A.5.
4.2 LLMFP Performance
We evaluate LLMFP on 9 tasks with two models: GPT-4o (gpt, ) and Claude 3.5 Sonnet (cla, ) with temperature 0. Each task comes with a natural language task description, background information on the scenarios or info collection API, and natural language queries. For single-step decision-making problems, expected output formats are also provided. With no task-specific example, LLMFP takes these inputs and delivers plans as outputs. We use optimal rate as the evaluation metric, that is, whether the plans are optimal under the current task scenario and query.
Baselines We compare LLMFP against 1) Direct: LLM direct plan generation, 2) CoT: chain-of-thought prompting (Wei et al., 2022) by asking LLMs to reason before generating the final answer, and 3) Code: prompts LLM to generate Python codes to solve the problem, allowing the use of any package or solver. For all baselines, we use both GPT-4o and Claude 3.5 Sonnet and also include a direct plan generation baseline with OpenAI o1-preview (o1p, ). All baselines are zero-shot with no task-specific examples. All baselines have the same input information as LLMFP, including task description, task background information or info collection API, and query. Please refer to Sec. A.7 for prompts of baselines.
Table 1: Optimal rate (%) comparison of LLMFP with baselines on 5 multi-constraint problems.
Method Coffee Workforce Facility Task Allocation Warehouse Average
[…]
Table 2: Optimal rate (%) comparison of LLMFP with baselines on 4 multi-step problems.
Method Blocksworld Mystery Blocksworld Movie Gripper Average
[…]
Results and Analysis We include the optimal rate comparison of LLMFP and baselines on 5 multi-constraint problems and 4 multi-step problems in Table 1 and 2. There are three key takeaways:
[…]
Second, among baselines, Code works better for multi-constraint problems, while Direct and CoT work better for multi-step problems. This validates that the skills required for solving different tasks are different. For multi-constraint problems, as heavy calculations are required to test every possible solution, it is hard for LLMs to directly plan, even with the strongest o1-preview model. For multi-step problems, since Code tries to use a PDDL planner, which requires LLM to generate fixed-format PDDL domain and problem files, it almost always fails to generate and call them correctly. While it is easier for LLMs to directly devise plans as the preconditions and effects of each action are easier to reason about than calculations. This further proves that LLMFP can tackle problems that are fundamentally different because it uses a universal and formal approach for all tasks.
Third, for Direct and CoT, Mystery Blocksworld’s performance degrades largely compared to Blocksworld, though they are fundamentally same problems. Changing predicate and action names to illogical names makes LLMs hard to understand the problem and generate reasonable plans. However, LLMFP still can obtain an overall strong optimal rate of 77.7% and 98.0% on Mystery Blocksworld for GPT-4o and Claude 3.5 Sonnet. This shows LLMFP is robust to obfuscated problems, as it can encode the problem as long as the problem is clearly defined regardless of the names.
To summarize, the experiment results show that LLMFP is capable of solving all 9 tasks with strong performance and is robust to fundamentally different and obfuscated problems. We also show the performance of LLMFP across iterations, time and cost statistics, and failure analysis of LLMFP for different tasks in Appendix A.2,  A.3, and A.4.
4.3 Effectiveness of LLMFP Components
We then validate each component of LLMFP with ablation experiments on 9 tasks. We examine the effectiveness of Definer, Formulator, and Self Assess & Modification by removing these components from our framework one at a time and comparing with LLMFP. We do not remove Code Generator and Result Formatter because they are the necessary components of LLMFP to deliver outputs. We use GPT-4o as the LLMs and optimal rate as the evaluation metric.
Results and Analysis We include the optimal rate performance comparison of LLMFP and baselines on 9 problems in Table 3. From Table 3 there are two key takeaways:
First, removing any of the 3 components from LLMFP negatively affects the performance. For multi-constraint problems, removing Definer, Formulator, and Self Assess & Modification lowers the optimal rate by 15.4%, 22.2%, and 21.9%. For multi-step problems, removing Formulator and Self Assess & Modification reduces the optimal rate by 87.4% and 12.4%.
Second, for different problems, the most effective components are different. Coffee degrades the most for No Definer; Warehouse and all multi-step problems drop the most for No Formulator; and Workforce decreases the most for No Self Assess & Modification. This again validates the diversity of the 9 problems and how they require different efforts to be successfully solved. Thus, LLMFP is an overall framework that could aid the process of planning from all aspects.
Note that Warehouse with No Definer is the only entry that reaches a higher optimal rate than LLMFP. We notice that the main failure case for LLMFP in Warehouse is: LLM overwrites the provided API get_distance and provides 1 as the output during Code Generation. For Warehouse with No Definer, this happens less frequently. This kind of issue could be alleviated by adding non-task-specific modifications to our framework, such as emphasizing direct use of the provided API in prompt or implementing a checker that checks for overwriting of provided APIs.
To summarize, all three components in LLMFP are effective and could account for diverse problems by providing comprehensive aids to solve planning problems.
Table 3: Optimal rate (%) comparison when removing some key components of LLMFP on all 9 tasks. LLMs used are GPT-4o.
Domain No Definer No Formulator No Self Assess & Modification LLMFP
Coffee 8.6 56.4 55.3 64.7
Workforce 84.4 80.5 27.3 92.2
Facility 61.6 53.7 53.7 70.7
Task Allocation 74.0 92.0 96.0 96.0
Warehouse 90 2.0 54.0 72.0
Average 63.7 56.9 57.2 79.1
Blocksworld N/A 0.2 95.3 96.2
Mystery Blocksworld N/A 0.0 74.4 77.7
Movie N/A 0.0 66.7 100.0
Gripper N/A 0.0 64.0 76.0
Average N/A 0.1 75.1 87.5
4.4 LLMFP with Task-Specific Example
[…]
Table 4: Optimal rate (%) comparison of LLMFP and LLMFP with one task-specific example in Formulator on Coffee task. Sets represent different types of what-if questions. LLMs are GPT-4o.
Method Set 1 Set 2 Set 3 Set 4 Set 5 Set 6 Set 7 Average
LLMFP 58.3 70.9 11.8 42.4 80.0 83.3 81.5 61.2
[…]
5 Conclusion
To account for the challenge of the trade-off between flexibility and task complexity for existing LLM planning works, we observe that the core of many planning problems lies in optimization problems and propose a universal approach for LLMs to solve planning problems. We propose LLMFP, a general-purpose LLM-based planning framework that captures key information from planning problems and formally formulates and solves them as optimization problems, with no task-specific examples needed. We test LLMFP on 9 diverse planning tasks with two LLMs to prove LLMFP’s capability of achieving strong performance over fundamentally very different tasks and show the effectiveness of components in our framework.
Limitations One limitation of LLMFP is that it needs clear and detailed task descriptions and queries. It is hard for LLMFP to define the problems’ goals and constraints if the task description is ambiguous or missing some important information. Another limitation is that since LLMFP encodes the planning problems into optimization problems and solves them with optimization solvers, the capability of LLMFP depends on the strength of the solver. For massive databases with numerous feasible plans, it may take a longer time for the solver to search for the optimal plan. Ways to mitigate this is to introduce heuristics to prioritize a portion of the choices or to switch from solving optimization problems to satisfaction problems for planning tasks that do not require optimality.
Appendix A Appendix
A.1 Domains
We test on 9 planning problems, including 5 multi-constraint decision making tasks and 4 multi-step tasks (Li et al., 2023; Valmeekam et al., 2024; Stein & Koller, 2023):
  • Coffee Coffee company sources beans from three suppliers with fixed capacity, roasts them at two facilities into dark or light coffee, and ships the roasted coffee to three retail locations. The company aims to minimize the total shipping and roasting cost while fulfilling the demand at each retail location. There are 266 different queries of 7 types in the dataset.
  • Workforce Assign workers to shifts; each worker may or may not be available on a particular day. The goal is to minimize the total payments to workers while fulfilling the shift requirements for two weeks. There are 231 different queries of 5 types in the dataset.
  • Facility A company currently ships its product from 5 plants to 4 warehouses. It is considering closing some plants to reduce costs. The goal is to decide which plant(s) to close to minimize transportation and fixed costs. There are 165 different queries of 4 types in the dataset.
  • Task Allocation Given tasks and three robots skilled in different tasks, the goal is to assign tasks to robots to minimize finish time. The finish time counts when the last robot stops working. There are 50 different queries describing the number of different tasks. This task and its data and queries are created by us.
  • Warehouse The robots need to finish tasks by visiting stations that are capable of accomplishing corresponding tasks. The goal is to find the list of stations while minimizing the total distance traveled. There are 50 different queries that include the random-length list of tasks to finish. This task and its data and queries are created by us.
  • Blocksworld The robot has four actions: pickup, putdown, stack, and unstack. The goal is to stack the blocks in the scene from their initial setup to a specific order with minimum steps. There are 602 different queries that describe blocks’ initial conditions and goal states.
  • Mystery Blocksworld An obfuscated version of Blocksworld. The action and predicate names are replaced with names that logically make no sense. There are 602 different queries that describe objects’ initial conditions and goal states.
  • Movie The goal is to get the required snacks, watch the movie, and recover the movie and counter to the original state with minimum steps. There are 21 different queries that describe objects’ initial conditions and goal states.
  • Gripper There are robots and balls in different rooms. Each robot, with two grippers, can pick, drop, and move balls between rooms. The goal is to place balls in specific rooms with minimum steps. There are 25 different queries describing objects’ initial conditions and goal states.
The queries for Coffee, Workforce, and Facility are what-if questions that change or add constraints to the existing scenarios. The queries of the rest tasks are different task initial and goal conditions. Task inputs including example queries are given in Appendix A.5.
A.2 LLMFP Performance over Iterations
Fig. 3 shows the performance of LLMFP over 5 iterations. The key observation is: number of iterations of Self Assess & Midification stage enables LLMFP to further improve the optimal rates, although we can observe that LLMFP does not need extensive iterations to achieve an overall satisfying performance.
Figure 3: Optimal rates of models across LLMFP Iterations
A.3 LLMFP Time and Cost Statistics and Analysis
Table 56, and 7 shows the time and cost statistics of LLMFP for GPT-4o on 9 tasks.
We could observe that for both LLM querying time and solver running time, all stages of LLMFP requires reasonable runtime. The longest runtime is prompting Formulator for Movie, which takes 31.9 seconds because it contains 9 actions thus requiring longer representation formulation.
From cost statistics we could observe that the average cost per query for all 9 tasks is around 0.1 dollar, indicating LLMFP is not costly.
[…]
Domain Definer Formulator Solver Formatter Code Gen. Self Assess & Mod.
Coffee 5.6 10.8 17.1 0.1 14.2 11.3
Workforce 3.4 5.1 8.3 11.0 3.1 7.8
Facility 3.8 7.5 6.4 0.7 4.3 4.0
Task Allocation 8.6 23.8 5.2 0.2 6.5 5.9
Warehouse 3.9 3.1 6.2 0.2 4.1 3.3
Blocksworld N/A 21.0 14.6 0.6 1.9 3.4
Mys. Blocksworld N/A 24.3 14.6 0.6 2.3 4.1
Movie N/A 31.9 15.6 0.5 11.8 6.5
Gripper N/A 18.3 16.0 6.9 11.2 7.0
[…]
Coffee Workforce Facility Task Allocation Warehouse
cost/query 0.139 0.140 0.083 0.081 0.085
[…]
Blocksworld Mystery Blocksworld Movie Gripper
cost/query 0.122 0.105 0.131 0.128
A.4 LLMFP Failure Case Analysis
Here we analyze the major failure cases for all 9 tasks.
A.4.1 Coffee
There are two major failure cases for Coffee tasks:
First, some queries are not clearly presented, indicating ambiguous information. Queries of Coffee tasks are what-if questions and are categorized into 7 sets. Each set is a type of question. We notice that the type “supply-roastery” asks queries like “What led to the decision to use supplier3 for the roasting facility at roastery1?”. To answer this question, it is both plausible to test “using supplier3” or to test “not using supplier3” to see the performance. However, the ground truth answer for this type of questions is to “not using supplier 3”. As confusing queries even for human, they are hard for LLMs to understand. Thus, for these queries, LLMFP sometimes generate codes with opposite meanings as what is expected.
Second, sometimes LLMFP Definer fails to consider all implicit constraints. The most easily neglectable implicit constraints are 1) the beans roasted in each roastery do not exceed the beans it receives, and 2) the beans ship from each roastery do not exceed the coffee it roasts. When any of the two constraints are missing, to minimize the cost, the model will automatically set the shipped beans or roasted coffee to be 0, assuming the company delivers coffee without sourcing beans or roasting coffee.
A.4.2 Workforce
There are two major failure cases for Coffee tasks:
First, sometimes LLMFP fails to understand the queries. Some of the queries asks questions like ‘Can Gu transition from Sun14 to Sun7 for work purposes?’. The meaning is to force Gu to work on Sun7 and take rest on Sun14. However, sometimes LLMFP builds variables to test both taking and not taking this transition, and returns solutions with less costs.
Second, sometimes when the solution space is large, it is hard to find the optimal solution within maximum runtime set for solver. We set the maximum solver runtime to be 15 minutes, which is exceeded when solving some hard queries.
A.4.3 Facility
Similarly as the first failure case of Coffee, some queries are not clearly presented. The queries are like “What justifies the opening of plant 0?”, which is confusing even for humans. Both opening plant 0 and closing plant 0 to report the costs make sense to answer this query. However, the ground truth meaning of this query is to close plant 0.
A.4.4 Task Allocation
LLMFP only fails one query in Task Allocation. The reason is the Formulator generates wrong values for robot finish time.
A.4.5 Warehouse
The major failure case for Warehouse is Code Generator overwrites the provided API get_distance and provide 1 as the output during Code Generation. Thus, the distance between each station is mistakenly set to be 1.
A.4.6 Blocksworld
One major failure case for Blocksworld is Code Generator fails to initialize the states of predicates correctly and thoroughly. Since the query will only meantion the predicates that are true, for example, block 1 is on block 2, but when initializing, LLMFP needs to initialize both mentioned states but also unmentioned states that are false. For example, block 2 is not on block 1. However, Code Generator sometimes fails to consider all unmentioned states.
A.4.7 Mystery Blocksworld
Similarly as Blocksworld, Mystery Blocksworld has same failure case. For Mystery Blocksworld, since the predicate and action names are not meaningful, more this kind of errors are made by GPT-4o. However, Claude seems to have better reasoning capability to support it from making more these errors.
A.4.8 Movie
There is no failure case for Movie.
A.4.9 Gripper
The major failure case for Gripper is when the solver fails to find the solution because there are some code generation errors, the Self Assess & Modification sometimes would think it is because the timestep is not enough, thus adding another loop within the original loop. However, this would result in the program to execute forever.
A.5 Inputs on 9 tasks
We include the inputs, which includes task description, background information or API, and example queries, for all 9 tasks in Fig. 4 - Fig. 12:
Figure 4: Task description, background information or API, and example queries for Coffee
Figure 5: Task description, background information or API, and example queries for Workforce
Figure 6: Task description, background information or API, and example queries for Facility
Figure 7: Task description, background information or API, and example queries for Task Allocation
Figure 8: Task description, background information or API, and example queries for Warehouse
Figure 9: Task description, background information or API, and example queries for Blocksworld
Figure 10: Task description, background information or API, and example queries for Mystery Blocksworld
Figure 11: Task description, background information or API, and example queries for Movie
Figure 12: Task description, background information or API, and example queries for Gripper
A.6 Example Outputs on Coffee tasks
In this section, we include the outputs of all models for the query “What is the potential impact of a 29% increase in demand at cafe cafe2?”. The optimal solution has total cost to be 2612.
A.6.1 Baselines
We include the outputs of four baselines for Coffee task in Fig. 13 - Fig. 15, and analyze the failure reason in the figure descriptions:
[…]
A.6.2 LLMFP
We include the outputs of all stages of LLMFP for Coffee task in Fig. 16 - Fig. 20:
[…]
A.7 Prompts
A.7.1 Baseline Prompt
Direct Prompt
You have a domain and a query under this domain that you need to fulfill.
The domain is: {task}
Query: {question}
You have the access to {info_api}
What is the plan to achieve my goal? Answer by fill in this JSON response directly with no explanation:
{output_format}
CoT Prompt
You have a domain and a query under this domain that you need to fulfill.
The domain is: {task}
Query: {question}
You have the access to {info_api}
What is the plan to achieve my goal? Lets think step by step, first reason about the problem and how to solve it, then answer by fill in the JSON:
Reason:
JSON response:
{output_format}
Code Prompt
You have a domain and a query under this domain that you need to fulfill.
The domain is: {task}
Query: {question}
You have the access to {info_api}
Please write Python code to help me find the plan to achieve my goal. You can import any package and use any solver.
At the end, save your found plan in a variable named feedback with the following format:
{output_format}
Please respond with code only and wrap your answer with ‘‘‘python and ‘‘‘:
A.7.2 LLMFP Prompt
We use general templates for all tasks. The full prompts for all tasks will be released soon. Here we show the templates we have for GPT-4o. Since Claude naturally considers more constraints and is more strict in assessing, we edit the prompts a little to account for the different traits of Claude, and prompts are also included in the codes.
We include the prompt template we use for GPT-4o as below: Definer Prompt
You are given a task description in natural language, and you want solve it by building an optimization problem for this task.
The task is: {task_description}
You have the access to {info_api}
To get started of building the optimization problem, what is the goal, decision variables, and constraints to consider for this task?
Specifically, consider:
Goal: define the objective trying to optimize
Decision variables: identify all the decision variables involved in the problem
Constraints: key requirement for decision variables; For every pair of decision variables, carefully consider relations (explicit, implicit, underlying assumption, unmentioned commonsense) between them and explicitly include as constraint to ensure all variables are connected with each other
Response with [[GOAL: ]], [[Decision Variables: ]], [[Constraints Reasoning: ]], and [[Constraints: ]] only with no explanation and no math formulas. Try to be thorough and include all needed information as much as you can.
Formulator Prompt for single-step multi-constraint problems
You are given a Query under a task description in natural language, and you want solve it by building an optimization problem for this task. You already have considered the goal and constraints of this optimization problem. Your job is, given access to existing variables or APIs and a specific natural language query, think about other variables needed to encode and solve this problem with Z3 SMT solver and describe the important attributes of variables as a JSON format description. Here are some example task-output pairs to refer to:
Example task 1: There are blocks of different colors and scores in the scene. You need to select required number of non-repeat blocks with required color, while maximizing the score.
Query: I previously want to select 20 blocks that are black or red, but now my demand raises 9%.
GOAL: Maximize the total score of selected blocks.
Decision Variables: Indexes of blocks selected
Constraint: The required number of selected blocks is met.
Constraint: The selected blocks are non-repeat.
Constraint: The selected blocks have required color.
Variable or API:
You have the access to function math.ceil() to round UP float to int and math.floor() to round DOWN float to int. Please ONLY use these to convert from float to int.
You have access to a BlockSearch API. BlockSearch.run(color:list) gives 1.all possible block ids of color in "color" list and 2.corresponding score info. BlockSearch.get_info(score_info, block_index) gives the score of certain block. ]
JSON description:
"variable_1": {
"name": "blocks",
"SMT_variable": true,
"number_of_variables": math.ceil(20 * 1.09),
"data_source": "BlockSearch.run()",
"value": "selecting math.ceil(20 * 1.09) blocks from black and red blocks",
"specific_requirement": "selected blocks are black or red; non-repeat blocks"
"variable_2": {
"name": "score",
"SMT_variable": true,
"number_of_variables": math.ceil(20 * 1.09),
"data_source": "BlockSearch.get_info()",
"value": "depends on variable_1 variables",
"specific_requirement": null
"variable_3": {
"name": "total_score",
"SMT_variable": true,
"number_of_variables": 1,
"data_source": "variable_2 variables",
"value": "sum of variable_2 variables",
"specific_requirement": "equal to sum of variable_2 variables, maximize"
Example task 2: Given a list of cities, you need to start from an origin city, non-repeatly visit each other city exactly once, and traval back to origin city, with minimized total distance travelled.
Query: Total number of cities is 10.
GOAL: Minimize the total travel distance.
Decision Variables: List of visited city indexes
Constraint: Start from and end with same city.
Constraint: Each city is visited exactly once and non-repeat.
Variable or API: You have access to a DistanceSearch() API. DistanceSearch.run() takes no argument and gives the distance info between cities, and DistanceSerarch.get_info(distance_info, city_1, city_2) gives the distance(a real number) between two cities.
Based on below examples, your task is to generate a JSON description to describe the problem.
JSON description:
"variable_1": {
"name": "cities",
"SMT_variable": true,
"number_of_variables": 10,
"how_to_pick": "selecting 10 cities from 10 cities",
"data_source": null,
"specific_requirement": "non-repeat cities"
"variable_2": {
"name": "distance",
"SMT_variable": true,
"number_of_variables": 10,
"how_to_pick": "depends on constraint_1 variables",
"data_source": "DistanceSearch.run(), DistanceSerarch.get_info()",
"specific_requirement": "distance between each city pair, and the distance back to origin city"
"variable_3": {
"name": "total_distance",
"SMT_variable": true,
"number_of_variables": 1,
"data_source": "variable_2 variables",
"value": "sum of variable_2 variables",
"specific_requirement": "equal to sum of variable_2 variables, minimize"
Now, based on the examples, solve the Query under new task setting and respond with similar format, please explicitly specify the action/requirement needed to fulfill query, and explicitly take into consideration every constraint mentioned:
The task is: {task}
Query: {question}
{definer_response}
Variable or API:
{info_api}
Think about variables needed to encode all constraints and goal, describe all important attributes of variables as a JSON format description.
Make sure to explicitly consider and include requirements/constraints needed to answer the query. Note that to answer the query "Why do xxx", you need to examine the effect of "not doing xxx" to provide reasons; and to answer the query "Why not do xxx", you need to examine the effect of "do xxx" to provide reasons.
Response with JSON only with no explanation.
Formulator Prompt for multi-step problems
You are given a Query under a task description in natural language, and you want solve it by building an optimization problem for this task. Your job is, given access APIs and a specific natural language query, think about variables needed to encode and solve this problem with Z3 SMT solver and describe the important attributes of variables as a JSON format description. Here is an example task-output pairs to refer to:
Example task:
You have to plan logistics to transport packages within cities via trucks and between cities via airplanes. Locations within a city are directly connected (trucks can move between any two such locations), and so are the cities. In each city there is exactly one truck and each city has one location that serves as an airport.
Here are the actions that can be performed and its preconditions and effects:
Load truck: Load a {package} into a {truck} at a {location} only if the package and the truck are both at location. After the Load truck action, the package is not at the location and is in the truck.
Load airplane: Load a {package} into an {airplane} at a {location} only if the package and the airplane are both at location. After the Load airplane action, the package is not at the location and is in the airplane.
Unload truck: Unload a {package} from a {truck} at a {location} only if the truck is at location and the package is in truck. After the Unload truck action, the package is not in the truck and is at the location.
Unload airplane: Unload a {package} from an {airplane} at a {location} only if the airplane is at location and the package is in airplane. After the Unload airplane action, the package is not in the airplane and is at the location.
Drive truck: Drive a truck from one {location_1} to another {location_2} within a {city} only if the truck is at location_1 and both location_1 and location_2 are both in city. After the Drive truck action, the truck is not at location_1 and is at location_2.
Fly airplane: Fly an airplane from one {location_1} in a city to another {location_2} in another city only if both locations are airport and the airplane is at location_1. After the Fly airplane action, the airplane is not at location_1 and is at location_2.
Query: You have 2 airplanes a0 and a1, 2 trucks t0 and t1, 2 cities c0 and c1, city c0 has location l0-0 and l0-0 is airport, city c1 has location l0-1 and l0-1 is airport, and a package p0. Initially, t0 is at location l0-0, t1 is at location l1-0, p0 is at location l1-0, a0 and a1 are at l0-0. The goal is to have p0 to be at l0-0.
API: You can assume you already know T as the input. You have access to a update_data() API that helps to update the predicate variables.
JSON description:
"objects": {
"variable_1": {
"name": "objects",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "query",
"value": "a dictionary that summarizes all objects in the problem: key package’, value [’p0’]; key airplane’, value [’a0’, a1’]; key truck’, value [’t0’, t1’]; key city’, value [’c0’, c1’]; key location’, value [’l0-0’, l0-1’]; key airport’, value [’l0-0’, l0-1’]",
"specific_requirement": null
"predicates": {
"variable_2": {
"name": "at",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "query, variable_1",
"value": "a dictionary of boolean variables representing whether an object is at a location at timestep: keys are (package/truck/airplane, location, timestep)",
"specific_requirement": "add constraint to initialize timestep 0 according to query, for unmentioned objects explicitly set it to be False"
"variable_3": {
"name": "in",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "query, variable_1",
"value": "a dictionary of boolean variables representing whether an object is in airplane or in truck: keys are [package, airplane/truck, timestep]",
"specific_requirement": "add constraint to initialize all values to be False at timestep 0"
"variable_4": {
"name": "in-city",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "query, variable_1",
"value": "a dictionary of boolean variables representing whether an location is in a city: keys are [location, city, timestep]",
"specific_requirement": "add constraint to initialize timestep 0 according to query, for unmentioned objects explicitly set it to be False"
"actions": {
"variable_5": {
"name": "load_truck",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "variable_1",
"value": "a dictionary of boolean variables representing whether load_truck action is performed for package, truck, location: keys are [package, truck, location, timestep]",
"specific_requirement": null
"variable_6": {
"name": "load_airplane",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "variable_1",
"value": "a dictionary of boolean variables representing whether load_airplane action is performed for package, airplane, location: keys are [package, airplane, location, timestep]",
"specific_requirement": null
"variable_7": {
"name": "unload_truck",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "variable_1",
"value": "a dictionary of boolean variables representing whether unload_truck action is performed for package, truck, location: keys are [package, truck, location, timestep]",
"specific_requirement": null
"variable_8": {
"name": "unload_airplane",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "variable_1",
"value": "a dictionary of boolean variables representing whether unload_airplane action is performed for package, airplane, location: keys are [package, airplane, location, timestep]",
"specific_requirement": null
"variable_9": {
"name": "drive_truck",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "variable_1",
"value": "a dictionary of boolean variables representing whether drive_truck action is performed for truck, location_from, location_to, city: keys are [truck, location, location, city, timestep]",
"specific_requirement": null
"variable_10": {
"name": "fly_airplane",
"SMT_variable": false,
"number_of_variables": 1,
"data_source": "variable_1",
"value": "a dictionary of boolean variables representing whether fly_airplane action is performed for airplane, location_from, location_to: keys are [airplane, location, location, timestep]",
"specific_requirement": null
"update": {
"step_1": {
"name": "action load_truck precondition and effect",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "query, variable_1, variable_5",
"value": "add constraints for preconditions and effects of load_truck",
"specific_requirement": "for each timestep t until T, for all package, truck, and location, assert that load_truck[package, truck, location, t] implies at[truck, location, t], at[package, location, t], not at[package, location, t+1], in[package, truck, t+1]"
"step_2": {
"name": "action load_airplane precondition and effect",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "query, variable_1, variable_6",
"value": "add constraints for preconditions and effects of load_airplane",
"specific_requirement": "for each timestep t until T, for all package, airplane, and location, assert that load_airplane[package, airplane, location, t] implies at[airplane, location, t], at[package, location, t], not at[package, location, t+1], in[package, airplane, t+1]"
"step_3": {
"name": "action unload_truck precondition and effect",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "query, variable_1, variable_7",
"value": "add constraints for preconditions and effects of unload_truck",
"specific_requirement": "for each timestep t until T, for all package, truck, and location, assert that unload_truck[package, truck, location, t] implies at[truck, location, t], in[package, truck, t], not in[package, truck, t+1], at[package, location, t+1]"
"step_4": {
"name": "action unload_airplane precondition and effect",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "query, variable_1, variable_8",
"value": "add constraints for preconditions and effects of unload_airplane",
"specific_requirement": "for each timestep t until T, for all package, airplane, and location, assert that unload_airplane[package, airplane, location, t] implies at[airplane, location, t], in[package, airplane, t], not in[package, airplane, t+1], at[package, location, t+1]"
"step_5": {
"name": "action drive_truck precondition and effect",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "query, variable_1, variable_9",
"value": "add constraints for preconditions and effects of drive_truck",
"specific_requirement": "for each timestep t until T, for all truck, location_from, location_to, city, assert that drive_truck[truck, location_from, location_to, city, t] implies at[truck, location_from, t], not at[truck, location_from, t+1], at[truck, location_to, t+1]"
"step_6": {
"name": "action fly_airplane precondition and effect",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "query, variable_1, variable_10",
"value": "add constraints for preconditions and effects of fly_airplane",
"specific_requirement": "for each timestep t until T, for all airplane, location_from, location_to, assert that fly_airplane[airplane, location_from, location_to, t] implies at[airplane, location_from, t], not at[airplane, location_from, t+1], at[airplane, location_to, t+1]"
"step_7": {
"name": "all_actions",
"SMT_variable": false,
"number_of_variables": "list of all actions",
"data_source": "variable_1, variable_5, variable_6, variable_7, variable_8, variable_9, variable_10",
"value": "for each timestep t until T, a list of all possible actions corresponding to different objects",
"specific_requirement": "for each timestep t until T, explicitly assert ONLY ONE action per timestep"
"step_8": {
"name": "unchanged predicate variables update",
"SMT_variable": null,
"number_of_variables": null,
"data_source": "update_data()",
"value": "update at, in, in-city using update_data()",
"specific_requirement": "update data with update_data()"
"goal": {
"step_9": {
"name": null,
"SMT_variable": null,
"number_of_variables": null,
"data_source": null,
"value": null,
"specific_requirement": "assert for timestep T, package p0 is at location l0-0"
Now, based on the example, solve the Query under new task setting and respond with similar format, please explicitly specify the action/requirement needed to fulfill query in your response:
The task is:
{task}
Query:
{question}
API: You have access to T as the input, so do NOT re-initialize T anywhere. You have access to a update_data(solver) API that helps to update the unchanged predicate variables. Please ONLY use this API to update unchaged predicates.
Response with JSON only with no explanation.
JSON description:
Code Generator Prompt
You are given a task description in natural language, a specific natural language query, available APIs and variables, and a JSON description that summarizes important variables that guide you to encode and solve the problem with SMT solver.
Your task is to generate steps and corresponding Python codes that utilizes Z3 SMT solver to solve the problem.
For the variables summarized in the JSON description:
(1) name represents the name of the variable
(2) SMT_variable indicates whether you should assign it as a normal variable or an SMT variable
SMT_variable Example: price = Int(’price’)
[…]
pick_ball = Bool(’pick ball’) # Boolean SMT variable
Normal variable Example: price = 100
flight_index = [1,2,3]
(3) number_of_variable represents the length of the variable
(4) data_source is the source for the variable to get the data
(5) value is, after you get needed data from any source, how you should assign these data to the variable
(6) specific_requirement is if there are any specific requirements that needs to be considered.
For the below problem, can you generate steps and corresponding Python codes to encode it? Do not include any explanations. You do not need to solve the problem or print the solutions.
The task is: {task}
Query: {question}
{definer_response}
Variable or API:
{info_api}
JSON variable representation:
{formulator_response}
Please use a SMT variable named total_cost when calculating the total cost. Please put the optimization goal at the end after all needed calculation and constraints additions.
Make sure your code add constraints to solver that considers and could answer the query. Note that to answer the query "Why do xxx", you need to examine the effect of "not doing xxx" to provide reasons; and to answer the query "Why not do xxx", you need to examine the effect of "do xxx" to provide reasons.
Initialize a Z3 optimizer solver = Optimize() at the beginning of the code.
Response with Python code only with no explanation.
Formatter Prompt
You are given a task description in natural language, a specific natural language query, pre-defined variables, and an execution feedback by running a Python Code that tries to solve the task.
The task is: {task}
Query: {question}
Execution feedback: {feedback}
Variable or API:
{info_api}
If the execution feedback is runtime errors, please return RUNTIME ERROR for JSON: and NULL for Correctness reasoning:.
If the execution feedback is cannot find the solution, please return CANNOT FIND SOLUTION for JSON and NULL for Correctness reasoning:.
If the execution feedback is not runtime errors, the execution feedback is the solved solution for this task. Only using the information from Execution feedback (do not use predefined variables), transform the execution feedback into a JSON format task plan by filling in the JSON below:
{output_format}
In addition, for Correctness reasoning, please explicitly answer one by one does the task plan satisfy these constraints? Include one sentece explanation for each constaint:
{{{definer_response}}}
Then explicitly answer and explain in one sentence: Does the task plan make sense and achievable in reality and meet commonsense?:
Please include your response here with no explanations:
JSON:
Correctness reasoning:
Self Assess & Modification Prompt
You are given a task and steps that tries to solve it as an optimization problem. The steps include:
1) specifying the goal and constraints of the optimization problem.
2) a JSON description that summarizes important variables that guide to encode and solve the problem with Z3 SMT solver.
3) the Python code to encode and solve the problem with Z3 SMT solver.
Your goal is to, based on the task description, specific query, available API or variables, and runtime execution feedback (it could either be an execution error or a generated plan if theres no runtime error), assess whether any steps 1-3 are correct.
The task is: {task}
Query: {question}
Variable or API:
{info_api}
Steps to judge:
1) {definer_response}
2) {formulator_response}
3) {code_generator_response}
Execution feedback: {feedback}
Based on the previous information, evaluate whether steps 1-3 are correct:
For Step 1: Does the step consider correct goal and all needed constraints? Are there unnecessary or missing constraints? Does the execution result make sense and achievable in reality and meet commonsense?
For Step 2: Do the variables explicitly consider the query? Do the variables explicitly consider and encode all provided constraints and goal? Does the order make sense? Does the execution result make sense and achievable in reality and meet commonsense?
For Step 3: Does the code create all needed variables? Does the code make up any ungiven information? Does the code correctly address the query? Does the execution result make sense and achievable in reality and meet commonsense?
Please reason the correctness with task context, rate each step with a binary score: 1 is correct, 0 is incorrect, think about how to modify in detail according to task and query, and modify the step if you think it is incorrect.
For Step 2 modification, please write in JSON format. For Step 3 modification, please write in Python and do noy change the content after line if solver.check() == sat: ’.
Your response format should be below, put NULL to How to mofify Reasoning and Modified Step if you think the step is correct, do not include extra explanation:
[[Step 1:
Correctness Reasoning:
Rating:
How to mofify Reasoning:
Modified Step 1(no explanation):
END
[[Step 2:
Correctness Reasoning:
Rating:
How to mofify Reasoning:
Modified Step 2(no explanation):
END
[[Step 3:
Correctness Reasoning:
Rating:
How to mofify Reasoning:
Modified Step 3(no explanation):
END
A.7.3 Prompts and output for LLMFP with MILP solver for Coffee Example
Definer and Formatter prompt remain exactly the same as for SMT solver.
We include the comparison of prompts for Formulator, Code Generator, and Self Assess & Modification in Fig. 21 to 23 and labelled all the differences with red. We then include the output of Formulator and Code Generator in Fig. 24 and Fig. 25. The key takeaway is it is very easy to switch from one solver to another with LLMFP, as the inner logic is same: building an optimization problem.
Figure 21: The Formulator prompt difference when switching from using Z3 SMT solver to Gurobi MILP solver.
Figure 22: The Code Genetator prompt difference when switching from using Z3 SMT solver to Gurobi MILP solver.
Figure 23: The Self Assess & Modification prompt difference when switching from using Z3 SMT solver to Gurobi MILP solver.
Figure 24: Output of Formulator for Coffee after switching the solver to MILP, using GPT-4o.
Figure 25: Output of Code Generator for Coffee after switching the solver to MILP, using GPT-4o.